OpenStudy (anonymous):

Let f(x)=18−7x−x^2. Find the open intervals on which f is increasing (decreasing). 1. f is increasing on the intervals 2. f is decreasing on the intervals

OpenStudy (anonymous):

is this calculus class or algebra?

OpenStudy (anonymous):

if it is algebra, then find the first coordinate of the vertex actually if it is calculus, do the same thing

OpenStudy (anonymous):

vertex is \(-\frac{b}{2a}\) which is your case is \(-\frac{-7}{2(-1)}=-\frac{7}{2}\)

OpenStudy (anonymous):

its for calculus

OpenStudy (anonymous):

ok well it makes no difference this is a parabola that opens down. it is increasing on \((-\infty, -\frac{7}{2})\) and decreasing elsewhere

OpenStudy (anonymous):

okay thank you